# An Empirical Study of Scaling Law for OCR

### 29 Dec 2023 v1

@@ Abstract
The laws of model size, data volume, computation and model performance have been extensively studied in the field of Natural Language Processing (NLP). However, the scaling laws in Optical Character Recognition (OCR) have not yet been investigated. To address this, we conducted comprehensive studies that involved examining the correlations between performance and the scale of models, data volume and computation in the field of text recognition. Conclusively, the study demonstrates smooth power laws between performance and model size, as well as training data volume, when other influencing factors are held constant. Additionally, we have constructed a large-scale dataset called REBU-Syn, which comprises 6 million real samples and 18 million synthetic samples. Based on our scaling law and new dataset, we have successfully trained a scene text recognition model, achieving a new state-of-the-art on 6 common test benchmarks with a top-1 average accuracy of 97.42.
@@ 3.1 Model Scaling
PARSeq [8] follows an encoder-decoder architecture. The encoder part utilizes the Vision Transformer (ViT) model to extract image features, while the decoder follows the same architecture as the pre-LayerNorm [83] Transformer decoder in this study utilizes twice the number of attention heads, where nhead = dmodel/32. In contrast to the standard ViT, the encoder removes the token and inputs all the output tokens into the decoder.
@@ Appendix D OCR Enhanced LMM
In the realm of large-scale models, we observe a distinct bifurcation into two primary categories: Large Language Models (LLMs) and Large Multimodal Models(LMMs). It is crucial to acknowledge that while LLMs are devoid of a visual component, LMMs’ visual branches demonstrate room for enhancement in terms of OCR capabilities[73]. This observation underscores the relative underdevelopment of OCR proficiency within large-scale models. Optical Character Recognition (OCR) tasks, however, offer a promising avenue to address this shortfall, thereby motivating our investigation into the benefits of integrating OCR with these models.
@@
Result and Analyze We performed a detailed comparative analysis to assess the accuracy of the QWen-VL-chat model, examining its performance with and without OCR integration, as delineated in Table 5. Our results reveal a significant improvement in the accuracy of the model for scene-based VQA tasks upon the integration of OCR. Additionally, there is a noticeable enhancement in document-based VQA tasks. These findings suggest that the incorporation of OCR not only enhances the model’s accuracy but also extends its generalization capabilities across diverse VQA scenarios. This evidence distinctly highlights the vital role that OCR inputs play in augmenting the performance of LVLM for downstream tasks. Furthermore, the improved accuracy with OCR integration underscores the model’s enhanced ability to interpret and analyze combined visual and textual data, thereby validating the efficacy of multimodal approaches in tackling complex analytical challenges.
VQA Visulization Analysis Our visual analysis of QWen-VL-Chat, with and without the OCR module, across varied datasets offers critical insights. In natural scene Visual Question Answering (VQA) tasks, QWen-VL-Chat encounters difficulties in detecting small text in images. The upper left corner of Fig 2, the model overlooks pertinent content, erroneously indicating its absence. Moreover, its tendency to inaccurately complete blurred text stems from its sophisticated linguistic abilities. This is evident in the upper right corner of Fig 2, where ’dr’ in ’45th parallel dr’ is incorrectly expanded to ’drive’. Notably, the model’s text misidentification issues, such as converting ’honghe’ to ’Hongte’ on a cigarette pack as depicted in the lower left corner of Fig 2 (mistaking the second ’h’ for a ’t’), are significantly mitigated by integrating our OCR module.

### 02 Jan 2024 v2

# Large OCR Model: An Empirical Study of Scaling Law for OCR

@@ Abstract
The laws of model size, data volume, computation and model performance have been extensively studied in the field of Natural Language Processing (NLP). However, the scaling laws in Optical Character Recognition (OCR) have not yet been investigated. To address this, we conducted comprehensive studies that involved examining the correlations between performance and the scale of models, data volume and computation in the field of text recognition. Conclusively, the study demonstrates smooth power laws between performance and model size, as well as training data volume, when other influencing factors are held constant. Additionally, we have constructed a large-scale dataset called REBU-Syn, which comprises 6 million real samples and 18 million synthetic samples. Based on our scaling law and new dataset, we have successfully trained a scene text recognition model, achieving a new state-of-the-art on 6 common test benchmarks with a top-1 average accuracy of 97.42. The models and dataset are publicly available at large-ocr-model.github.io.
@@ 3.1 Model Scaling
PARSeq [8] follows an encoder-decoder architecture. The encoder part utilizes the Vision Transformer (ViT) model to extract image features, while the decoder follows the same architecture as the pre-LayerNorm [83]. Transformer decoder in this study utilizes twice the number of attention heads, where nhead = dmodel/32. In contrast to the standard ViT, the encoder removes the token and inputs all the output tokens into the decoder.
@@ Appendix D OCR Enhanced LMM
In the realm of large-scale models, we observe a distinct bifurcation into two primary categories: Large Language Models (LLMs) and Large Multimodal Models(LMMs). It is crucial to acknowledge that while LLMs are devoid of a visual component, LMMs’ visual branches demonstrate room for enhancement in terms of OCR capabilities[73]. This observation underscores the relative underdevelopment of OCR proficiency within large-scale models. However, OCR tasks offer a promising avenue to address this shortfall, thereby motivating our investigation into the benefits of integrating OCR with these models.
@@
Result and Analyze We performed a detailed comparative analysis to assess the accuracy of the QWen-VL-chat model, examining its performance with and without OCR integration, as delineated in Table 5. Our results reveal a significant improvement in the accuracy of the model for scene-based VQA tasks upon the integration of OCR. Additionally, there is a noticeable enhancement in document-based VQA tasks. These findings suggest that the incorporation of OCR not only enhances the model’s accuracy but also extends its generalization capabilities across diverse VQA scenarios. This evidence distinctly highlights the vital role that OCR inputs play in augmenting the performance of LMM for downstream tasks. Furthermore, the improved accuracy with OCR integration underscores the model’s enhanced ability to interpret and analyze combined visual and textual data, thereby validating the efficacy of multimodal approaches in tackling complex analytical challenges.
VQA Visulization Analysis Our visual analysis of QWen-VL-Chat, with and without the OCR module, across varied datasets offers critical insights. In natural scene VQA tasks, QWen-VL-Chat encounters difficulties in detecting small text in images. The upper left corner of Fig 2, the model overlooks pertinent content, erroneously indicating its absence. Moreover, its tendency to inaccurately complete blurred text stems from its sophisticated linguistic abilities. This is evident in the upper right corner of Fig 2, where ’dr’ in ’45th parallel dr’ is incorrectly expanded to ’drive’. Notably, the model’s text misidentification issues, such as converting ’honghe’ to ’Hongte’ on a cigarette pack as depicted in the lower left corner of Fig 2 (mistaking the second ’h’ for a ’t’), are significantly mitigated by integrating our OCR module.